OpenStudy (zenmo):

Find the limit:

OpenStudy (zenmo):

\[\lim_{t \rightarrow 0}\frac{ t^2-+t-2 }{ (t-1) }\]

OpenStudy (zenmo):

The bottom is an absolute value of (t-1). Wasn't sure how to do the absolute value sign in the equation format.

OpenStudy (zenmo):

oops, the top is \[t^2+t-2\]

OpenStudy (irishboy123):

just stuff in t = 0, see where you go. and sort out the denominator:-) you can say |x| = absolute value of x.

OpenStudy (anonymous):

\[\lim_{t \rightarrow 0}\frac{ t^2+t-2 }{ \left| t-1 \right| }\] \[when~ t \rightarrow 0,t-1<0,\] \[\lim_{t \rightarrow 0}\frac{ t^2+t-2 }{ -t+1 }=\frac{ 0+0-2 }{ -0+1 }=-2\]

OpenStudy (zenmo):

\[t -> 0, t-1 < 0,\] How come it is -t+1 instead of -(t+1) = -t - 1? Other then that, I understand the problem.

OpenStudy (anonymous):

if \[\left| x-a \right|<0,`then~\left| x=a \right|=-\left( x-a \right)=-x+a\]

OpenStudy (zenmo):

Oh, I made a slight error with the "+" and "-," I read it wrong. I understand completely now. THanks! :)

OpenStudy (anonymous):

correction \[\left| x-a \right|=-\left( x-a \right)=-x+a\]

OpenStudy (zenmo):

One more question, would we have to solve for t-1 > 0 next or just t-1 < 0?

OpenStudy (zenmo):

@surjithayer

OpenStudy (anonymous):

as t approaches 0 0<1 hence t<1 t-1<0 so in both ways t-1<0 i.e.,t approaches 0-and t approaches 0+

OpenStudy (zenmo):

@surjithayer To clarify, there are two answers: -2 & 2?

OpenStudy (irishboy123):

you meant to ask this, i think: \[\lim_{t \rightarrow 0}\frac{ t^2-+t-2 }{ |t-1| }\]? so if t = 0 ?????? you have what?